# Racing Thoughts: Explaining Large Language Model Contextualization Errors

### v1

@@ Abstract
The profound success of transformer-based language models can largely be attributed to their ability to integrate relevant contextual information from an input sequence in order to generate a response or complete a task. However, we know very little about the algorithms that a model employs to implement this capability, nor do we understand their failure modes. For example, given the prompt John is going fishing, so he walks over to the bank. Can he make an ATM transaction?, a model may incorrectly respond Yes if it has not properly contextualized bank as a geographical feature, rather than a financial institution. We propose the LLM Race Conditions Hypothesis as an explanation of contextualization errors of this form. This hypothesis identifies dependencies between tokens (e.g., bank must be properly contextualized before the final token, ?, integrates information from bank), and claims that contextualization errors are a result of violating these dependencies. Using a variety of techniques from mechanistic intepretability, we provide correlational and causal evidence in support of the hypothesis, and suggest inference-time interventions to address it.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated a remarkable capacity for accomplishing a wide variety of language generation and classification tasks (Team et al., 2023; Touvron et al., 2023; Brown et al., 2020). Many of these capabilities crucially rely on contextualization—integrating relevant contextual information from the input sequence at inference time. Indeed, LLMs’ profound success at contextualization has given rise to a novel method of adapting models to new tasks: in-context learning (Achiam et al., 2023; Wei et al., 2022). Although contextualization is fundamental to the success of large language models, we understand fairly little about the mechanisms underlying successful contextualization. At an implementation level, we understand that attention heads allow information to be read from and written to token representations across different positions (Elhage et al., 2021). However, we are currently in the dark about the algorithm-level factors that result in success or failure of contextualization.
Consider the contextualization error presented in Fig. 1 (left). This error was generated by a state of the art LLM, Gemini [^1] Example generated on 9/11/2024. (Team et al., 2023). Crucially, it betrays that the model has not properly contextualized the representation of bank. The model first interprets the word to refer to a geographic feature, and then interprets that same instance of bank in the same context as a financial institution.
@@
We verify the LLM Race Conditions Hypothesis using state-of-the-art techniques in mechanistic interpretability, offering qualitative and quantitative evidence as well as several inference-time interventions to ameliorate the problem. We conclude by arguing that race conditions are endemic to feedforward language models, and point toward potential solutions.
Figure 1: (Left) A contextualization error found in Gemini, a frontier LLM (Team et al., 2023). In this dialogue, the user’s first message implies that the correct word sense of bank is as a geographical feature (i.e., a river bank). Though the model recognizes this in its first reply, it fails to maintain this word sense of bank when probed in the very next user message, instead defaulting to the interpretation of bank as a financial institution. (Right) Illustrating the LLM Race Conditions Hypothesis. The Race Conditions Hypothesis suggests that contextualization errors result from out-of-order contextualization over the layers of an LLM. In this case, the question tokens are contextualizing with the polysemous word bank in an early layer, before its word sense is resolved via contextualizing with cue tokens.
@@
    We identify a simple and robust failure mode of LLM contextualization, which impacts state-of-the-art models across three distinct datasets.
@@ 2 Methods
For each subject entity and cue, we specify two questions, the correct answer to one of which is “yes” and the other is “no.” In the above example, the questions would be Is it a geographical feature? and Is it a financial institution?. Together, these questions comprise a question pair. We measure model performance in terms of question pairs: a question pair is marked correct only if the model answers both questions accurately. This controls for a model’s potential bias towards responding “yes” (Dentella et al., 2023).
For each subject entity and cue, we systematically vary the number of distractors included in the prompt, as well as the position of the cue amongst these distractors. To ensure that our analyses are robust to the particular distractors that we randomly select, we include three independent random samples of distractors for each subject entity and cue.
@@
We leverage the Patchscopes framework, which encompasses a set of techniques for inspecting the information that is encoded within a particular intermediate representation of a transformer (Ghandeharioun et al., 2024a). This intermediate or source representation is copied to a target position in a separate text prompt of the LLM, constructed such that the output will be informative regarding the content of the source. Using the terminology of Ghandeharioun et al., the source representation is defined by a tuple, $\bm{h}\equiv(S,i,\mathcal{M},l)$, where $S$ refers to a sequence of $n$ tokens, $i$ refers to a position in that sequence ($i\in[1,...,n]$), $\mathcal{M}$ refers to a model with $L$ layers, and $l$ refers to a particular layer in $\mathcal{M}$. In other words, a source representation is the result of running $\mathcal{M}$ on $S$, then extracting the residual stream state of position $i$ from layer $l$.
The Patchscopes target is defined as another tuple, $\bar{\bm{h}}\equiv(T,i^{*},\mathcal{M},l^{*})$, where $T$ is a target prompt consisting of $n^{*}$ tokens, $i^{*}$ refers to a position in $T$, $\mathcal{M}$ refers to the same model that generates the source representation, $\bm{h}$, and $l^{*}$ refers to a layer in $\mathcal{M}$. A patchscope intervention refers to replacing the representation $\bar{\bm{h}}$ with $\bm{h}$ at inference time. $T$ is constructed such that the output of $\mathcal{M}$ will be informative regarding the content of $\bm{h}$.
@@
We study gemma-2-9b-it (Team et al., 2024), and replicate using another commonly used open-weight model in § F. These models have been specifically tuned for instruction-following and chat capabilities and represent a sample of the state of the art in open-weight models.
@@ 3 Behavioral Failure Modes
In this section we document a simple and robust failure mode of contextualization in LLMs. Inspired by recent work investigating context integration in multi-document question answering (Liu et al., 2024) and long-context retrieval (Kamradt, 2024), we inject randomly-selected distractor sentences into prompts that contains a context-sensitive question. We find that the presence of distractors significantly deteriorates a model’s ability to integrate relevant contextual information when queried with context-dependent questions. For each of the three domains described in § 2, we generate six datasets, containing between zero and five distractors. We then evaluate the LLM’s ability to answer question pairs correctly in the face of distractors. The results are shown in Fig. 2(a). We observe that, across the board, distractor text greatly harms a model’s ability to properly contextualize.
(a) Model accuracy vs. number of distractors. We find that injecting distractor text into the prompt causes sharp performance degredations across all datasets.
(b) Performance on the 5-distractors dataset, disagreggated by the position of the cue amongst distractor sentences.
Figure 2: Behavioral results across all three datasets for Gemma-2. We find that injecting distractors routinely engenders contextualization errors.
Additionally, Liu et al. (2024) show serial position effects—multi-document question-answering performance is highest when the relevant document is presented at the beginning or at the end of a set of documents, and lowest when it occurs in the middle. However, their work studies contexts that contain thousands of tokens, whereas our longest questions are on the order of 100-200 tokens. Interestingly, we also observe sensitivity to the position of relevant information: Fig. 2(b) presents results from our five-distractor datasets, indexed by the position of the cue amongst the distractors. The model often performs better when the cue occurs before the distractors, and sometimes (i.e., in the Polysemous dataset) demonstrates the U-shaped curve found in Liu et al. (2024). See § F for a replication using another open-weight model.
@@ 4 The LLM Race Conditions Hypothesis
The LLM Race Conditions Hypothesis can be broken into two separate, but related, claims: (1) There exists a critical window in which the question tokens are contextualized with the subject entity (and other tokens). Any further refinement of the subject entity after the critical window will not be reflected in model responses. (2) Model failures are driven by unfinished contextualization of the subject entity at early layers (i.e., before the end of the critical window). In the following sections, we provide correlational and causal evidence in support of both of these claims.
For all of the analyses to follow, we select a dataset slice (i.e., number of distractors and cue position) that results in as close to 50% accuracy as possible for each model-dataset pair. This is helpful because some interventions have the goal of increasing accuracy, and others have the goal of decreasing accuracy through ablation. Note that 50% is not chance accuracy, as we are measuring in terms of question pairs, which require two correct answers (corresponding to one “yes” and one “no”). Thus, 25% is chance accuracy. See § A for details.
4.1 Evidence Supporting the Race Conditions Hypothesis: A Critical Window
We first verify the existence of a critical window, i.e., a set of layers in which the question token representations integrate information from the broader context, and after which they commit to an answer. We support this by (1) analyzing attention patterns, and finding that attention to the subject entity peaks in the middle layers, (2) performing a logit lens analysis of the final token of the prompt, and showing that the model’s ultimate decision is decipherable halfway through the network, and (3) performing a causal analysis by modulating attention patterns, and demonstrating that contextualization fails to impact downstream performance after the middle layers of the model.
@@ 4.1 Evidence Supporting the Race Conditions Hypothesis: A Critical Window
In each dataset, correct answers crucially rely on the contextualization of a subject entity (i.e., a polysemous word, a profession, or a country). We begin our investigation by directly analyzing attention patterns, a common first step in understanding how dependencies between tokens are represented in transformers (Clark et al., 2019; Htut et al., 2019; Lepori et al., 2024). We analyze the attention paid to the subject entities at each layer by all subsequent tokens (which comprise the question). Specifically, for each layer $l$, we compute the mean attention paid to the (final token of the) subject entity over all heads by all subsequent tokens, which we call the attention mass:
@@
$\text{diff}_{y,l}=\frac{\sum_{h_{l}\in D_{y}^{+}}\text{log-odds}(h_{l}))}{|D_{y}^{+}|}-\frac{\sum_{h_{l}\in D_{y}^{-}}\text{log-odds}(h_{l}))}{|D_{y}^{-}|}$ (2)
This value quantifies when the logit differences of a successful model generation differs from the logit differences of a failure, while controlling for intrinsic biases toward producing affirmative or negative answers that occur throughout the layers of a model (see § B). From Fig. 4, we see that there is no difference between success and failures until the middle layers of the network. After this point, there is clear signal indicating the answer that the model will ultimately generate. Thus, we might pin down the end of the critical window approximately to layer 20, or shortly thereafter. See § B for disaggregated details, which further support the conclusions presented here, and additional analysis.
@@
Figure 5: Attention Ablation Results for Gemma-2. We find that ablating cue tokens and distractor tokens both have the intended impact on performance – ablating cue tokens drops performance and ablating distractors increases performance. Most notably, however, we find that these interventions only impact model performance in the first half of layers.
Given the qualitative results obtained by analyzing attention and logit lens, we should expect that contextualization should only impact model performance in early and middle layers, before the question tokens finish reading from the subject token(s). We perform causal interventions on the attention maps produced throughout the transformer layers to test this. Specifically, we perform attention ablations: zeroing out the attention map entry that defines how much attention one token pays to another and renormalizing the resulting values. This effectively removes the influence of one token on another within the layers in which we apply this intervention. Formally, let $T_{E}$ be the set of tokens to edit and $T_{A}$ be the set of tokens to ablate. Attention ablations are defined as follows: $\forall_{e\in T_{E},a\in T_{A}}A_{e,a}\leftarrow 0$. We use the transformerlens package to implement this intervention (Nanda & Bloom, 2022).
We perform two types of attention ablations: distractor ablation, where $T_{A}$ corresponds to the tokens comprising distractor sentences, and cue ablation, where $T_{A}$ corresponds to tokens comprising the cue. Distractor ablation mimics the no distractors setting—up to a difference in positional embeddings—over the layers in which it is applied, and should boost performance. Cue ablation should leave the model unable to properly contextualize the subject entity, harming accuracy.
We perform distractor ablation and cue ablation in blocks of 5 layers, and analyze the impact of attention ablation over the course of the model. Our main Gemma-2 results are presented in Fig. 5, and a replication with another LLM is presented in § F. In all cases, we find that distractor ablation reliably increases performance in early layers, and cue ablation reliably decreases performance in early layers. Furthermore, contextualization interventions stop impacting model performance at approximately halfway through the model layers—almost precisely when logit lens reveals contextualization of the final question token.
4.2 Evidence Supporting the Race Conditions Hypothesis: Subject Entity Contextualization
Next, we verify that model failures are caused by unresolved contextualization of the subject entity in early layers of the model (i.e., before the end of the critical window). We support this by (1) analyzing the content of subject entity representations over layers, and finding that distractors deteriorate contextualization, (2) performing causal patching analyses, which show that contextualization of the subject entity mediates model performance and that well-contextualized representations emerge in late layers of the model.
@@ 4.2 Evidence Supporting the Race Conditions Hypothesis: Subject Entity Contextualization
First, we qualitatively investigate how the content of subject entity representations evolve over layers of processing. To do so, we employ an open-ended Patchscope (Ghandeharioun et al., 2024a) inspired by the SelfIE technique (Chen et al., 2024). This technique formulates a patchscope where the target prompt, $T$, is an interpretation prompt—a prompt designed to encourage the model to describe the hidden representation. Source representations are extracted from $i$, where $i$ is position of the last token of the subject entity, and $l$ ranges over every even layer (i.e., $l\in[0,2,4...,N]$) in an LLM, $\mathcal{M}$. The source representations are patched into a target defined by $T=$Tell me about X X X”, $i^{*}=$ the index of the X tokens, and $l^{*}=3$ (following Chen et al., 2024). We then condition the model to respond with “Sure! In this context, the word refers to” and allow the model to freely generate 15 tokens. This process ideally results in a natural language description of the content of the hidden representation.
We employ this technique to investigate how the word sense of the subject entities in our polysemous word dataset evolve over the layers of the model. Following prior work (Ghandeharioun et al., 2024b), we use another LLM [^2] PaLM2 text-unicorn@001 (Anil et al., 2023) to autoscore these generations as being indicative of one sense or another. See § C for details and § D for example interpretations. We present the results of this analysis in Fig. 6, showing cumulative accuracies over layers. A question pair is marked correct at layer $m$ if the model has properly contextualized the intermediate representation in at least one layer [0, $m$]. In Fig. 6 (left), we compare question pairs with and without distractors (using the same dataset partition described in § A). We find that the presence of distractors delays proper contextualization of intermediate representations, shifting the cumulative accuracy curve for question pairs with distractors to the right. This delay may limit the number of layers within the critical window where the subject tokens are well-contextualized. Furthermore, in Fig. 6 (right), we compare distractor question pairs that the model ultimately answers correctly with those that the model ultimately answers incorrectly. We find that cumulative accuracy is lower overall in failure cases, indicating that subject entities are sometimes never properly contextualized during model failures.
@@
At minimum, the LLM Race Conditions Hypothesis implies that improperly contextualized subject-entity representations somewhere in the model are causing incorrect responses. Behaviorally, we know that the model achieves ceiling performance without distractors, implying that the model contains a properly contextualized subject-entity representation. Can we boost performance in the presence of distractors by injecting subject-entity representations generated by models when they are faced with data without distractors? To test this, we perform a Patchscopes intervention. Source representations are specified where $S$ corresponds to a no-distractors prompt, $i$ corresponds to the position of the last token of the subject entity in $S$, $\mathcal{M}$ corresponds to an LLM with $N$ layers, and $l\in[0,2,4,6,...,M]$, where $M=N/2$. Target representations are specified where $T$ corresponds to a prompt with distractors, $i^{*}$ refers to the position of the last token of the subject entity in $T$, and $l^{*}=l$. We refer to this intervention as cross-patching. We search over layers for each question pair separately. If any source layer-target layer combination results in correct generations for both questions, we consider the question pair correct. From Fig. 7, one can see that cross-patching reliably boosts performance across all datasets. See § F for a replication using another open-weight LLM.
@@
The LLM Race Conditions Hypothesis requires that subject entity contextualization must happen either before or during the critical window of question contextualization. Otherwise, the properly contextualized subject entity would not be integrated in the final prediction. To test this, we employ a backpatching intervention described in Biran et al. (2024): source representations are extracted from a prompt, $S$, at some position $i$ and layer $l$. Target representations are defined by $T=S$, $i^{*}=i$ and $l^{*}<l$. In our case, we let $l^{*}\in[0,2,4,...,M]$ and $l\in[M,M+2,M+4,...,N]$, where the LLM $\mathcal{M}$ has $L$ layers and $M=L/2$. For each layer in the target set, we attempt to patch in a representation from each layer in the source set. In Fig. 7, we find that this reliably increases performance across all datasets. See § F for a replication using another open-weight LLM.
@@
In the previous experiment, the backpatched representation undergo further processing through the MLP and attention mechanisms after the target layer, $l^{*}$. Perhaps the performance advantages conferred by backpatching are a result of this extra computation, rather than the relative ordering of question contextualization and subject entity contextualization? In other words: what if the representations in the source layer, $l$, are not actually well-contextualized without further processing? If they were, perhaps the model would have answered correctly after all! To address this, we rerun the same experiment and source-target layer search, except we patch the source representation into every layer in between the source and target. Formally, if $l^{\prime}$ indicates the set of target layers, $l^{\prime}=[l^{*},l^{*}+1,...,l]$. This effectively freezes the representation, rendering it insensitive to MLP and attention mechanisms throughout $l^{\prime}$. From Fig. 7, we see that this setting produces similar result as the standard backpatching setting. Thus, the ordering of contextualization, rather than extra computation, engenders higher performance. See § E for further analysis, including results from every source-target layer combination. See § F for a replication using another open-weight LLM.

### v2

# Racing Thoughts: Explaining Contextualization Errors in Large Language Models

@@ Abstract
The profound success of transformer-based language models can largely be attributed to their ability to integrate relevant contextual information from an input sequence in order to generate a response or complete a task. However, we know very little about the algorithms that a model employs to implement this capability, nor do we understand their failure modes. For example, given the prompt John is going fishing, so he walks over to the bank. Can he make an ATM transaction?, a model may incorrectly respond Yes if it has not properly contextualized bank as a geographical feature, rather than a financial institution. We propose the LLM Race Conditions Hypothesis as an explanation of contextualization errors of this form. This hypothesis identifies dependencies between tokens (e.g., bank must be properly contextualized before the final token, ?, integrates information from bank) and claims that contextualization errors are a result of violating these dependencies. Using a variety of techniques from mechanistic intepretability, we provide correlational and causal evidence in support of the hypothesis and suggest inference-time interventions to address it.
Racing Thoughts: Explaining Contextualization Errors in Large Language Models
Michael A. Lepori Google DeepMind Brown University michael_lepori@brown.edu                        Michael C. Mozer & Asma Ghandeharioun Google DeepMind {mcmozer,aghandeharioun}@google.com
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated a remarkable capacity for accomplishing a wide variety of language generation and classification tasks (Team et al., 2023; Touvron et al., 2023; Brown et al., 2020). Many of these capabilities crucially rely on contextualization—integrating relevant contextual information from the input sequence at inference time. Indeed, LLMs’ profound success at contextualization has given rise to a novel method of adapting models to new tasks: in-context learning (Achiam et al., 2023; Wei et al., 2022). Although contextualization is fundamental to the success of large language models, we understand fairly little about the mechanisms underlying successful contextualization. At an implementation level, we understand that attention heads allow information to be read from and written to token representations across different positions (Elhage et al., 2021). However, we are currently in the dark about the algorithm-level factors that result in the success or failure of contextualization.
Consider the contextualization error presented in Fig. 1 (left). This error was generated by a state of the art LLM, Gemini [^1] Example generated on 9/11/2024. (Team et al., 2023). Crucially, it betrays that the model has not properly contextualized the representation of bank. The model first interprets the word to refer to a geographical feature, and then interprets that same instance of bank in the same context as a financial institution.
@@
We verify the LLM Race Conditions Hypothesis using state-of-the-art techniques in mechanistic interpretability, offering qualitative and quantitative evidence as well as several inference-time interventions to ameliorate the problem. We conclude by arguing that race conditions are endemic to feedforward language models and point toward potential solutions. [^2] Code and data are available here.
Figure 1: (Left) A contextualization error found in Gemini, a frontier LLM (Team et al., 2023). In this dialogue, the user’s first message implies that the correct word sense of bank is as a geographical feature (i.e., a river bank). Though the model recognizes this in its first reply, it fails to maintain this word sense of bank when probed in the very next user message, instead defaulting to the interpretation of bank as a financial institution. (Right) Illustrating the LLM Race Conditions Hypothesis. The Race Conditions Hypothesis suggests that contextualization errors result from out-of-order contextualization over the layers of an LLM. In this case, the question tokens are contextualized with the polysemous word bank in an early layer, before its word sense is resolved via contextualizing with cue tokens.
@@
    We identify a simple and robust failure mode of LLM contextualization, which impacts state-of-the-art models across three datasets.
@@ 2 Methods
For each subject entity and cue, we specify two questions – the correct answer to one of which is “yes” and the other is “no.” In the above example, the questions would be Is it a geographical feature? and Is it a financial institution?. Together, these questions comprise a question pair. We measure model performance in terms of question pairs: a question pair is marked correct only if the model answers both questions accurately. This controls for a model’s potential bias towards responding “yes” (Dentella et al., 2023).
For each subject entity and cue, we systematically vary the number of distractors included in the prompt, as well as the position of the cue amongst these distractors. To ensure that our analyses are robust to the particular distractors that we randomly select, we include three independent random samples of distractors for each subject entity and cue. Importantly, these datasets are not meant to be reflective of real-world contexts, but are instead constructed to stress-test models. By exacerbating problems that are present in normal LLM computation, we seek to gain a deeper understanding of contextualization within LLMs.
@@
We leverage the Patchscopes framework, which encompasses a set of techniques for inspecting the information that is encoded within a particular intermediate representation of a transformer (Ghandeharioun et al., 2024a). This intermediate or source representation is copied to a target position in a separate text prompt of the LLM, constructed such that the output will be informative regarding the content of the source. Using the terminology of Ghandeharioun et al., the source representation is defined by a tuple, $h\equiv(S,i,\mathcal{M},l)$, where $S$ refers to a sequence of $n$ tokens, $i$ refers to a position in that sequence ($i\in[1,...,n]$), $\mathcal{M}$ refers to a model with $L$ layers, and $l$ refers to a particular layer in $\mathcal{M}$. In other words, a source representation is the result of running $\mathcal{M}$ on $S$, then extracting the residual stream state of position $i$ from layer $l$.
The patchscopes target is defined as another tuple, $\bar{h}\equiv(T,i^{*},\mathcal{M},l^{*})$, where $T$ is a target prompt consisting of $n^{*}$ tokens, $i^{*}$ refers to a position in $T$, $\mathcal{M}$ refers to the same model that generates the source representation, $h$, and $l^{*}$ refers to a layer in $\mathcal{M}$. A patchscope intervention refers to replacing the representation $\bar{h}$ with $h$ at inference time. $T$ is constructed such that the output of $\mathcal{M}$ will be informative regarding the content of $h$.
@@
We primarily study gemma-2-9b-it (Team et al., 2024), and replicate our results using another similarly-sized open-weight model in § G. Finally, we replicate our results using a smaller model in the Gemma family, gemma-2-2b-it, in § H. These models have been specifically tuned for instruction-following and chat capabilities and represent a sample of the state of the art in open-weight models. This work required approximately 50 GPU-hours, using Nvidia A100 GPUs.
@@ 3 Behavioral Failure Modes
In this section we document a simple and robust failure mode of contextualization in LLMs. Inspired by recent work investigating context integration in multi-document question answering (Liu et al., 2024) and long-context retrieval (Kamradt, 2024), we inject randomly-selected distractor sentences into prompts that contains a context-sensitive question. We find that the presence of distractors significantly deteriorates a model’s ability to integrate relevant contextual information when queried with context-dependent questions. For each of the three domains described in § 2, we generate six datasets, containing between zero and five distractors. We then evaluate the LLM’s ability to answer question pairs correctly in the face of distractors. The results are shown in Fig. 2. We observe that, across the board, distractor text greatly harms a model’s ability to properly contextualize. Furthermore, we observe different performance depending on where the cue is interleaved within the distractor sentences (See § A).
Figure 2: Model accuracy vs. number of distractors. We find that injecting distractor text into the prompt causes sharp performance degradations across all datasets.
@@ 4 The LLM Race Conditions Hypothesis
The LLM Race Conditions Hypothesis can be broken into two separate, but related, claims:
(1) There exists a critical window in which the question tokens are contextualized with the subject entity (and other tokens). Any further refinement of the subject entity after the critical window will not be reflected in model responses. (2) Model failures are driven by unfinished contextualization of the subject entity at early layers (i.e., before the end of the critical window). In the following sections, we provide correlational and causal evidence in support of both of these claims.
For all of the analyses to follow, we select a dataset slice (i.e., number of distractors and cue position) that results in as close to 50% accuracy as possible for each model-dataset pair. This is helpful because some interventions have the goal of increasing accuracy, and others have the goal of decreasing accuracy through ablation. Note that 50% is not chance accuracy, as we are measuring in terms of question pairs, which require two correct answers (corresponding to one “yes” and one “no”). Thus, 25% is chance accuracy. See § B for details and examples from each dataset.
4.1 A Critical Window
We first verify the existence of a critical window, i.e., a set of layers in which the question token representations integrate information from the broader context, and after which they commit to an answer. We support this by (1) analyzing attention patterns and finding that attention to the subject entity peaks in the middle layers, (2) performing a logit lens analysis of the final token of the prompt and showing that the model’s ultimate decision is decipherable halfway through the network, and (3) performing a causal analysis by modulating attention patterns and demonstrating that contextualization fails to impact downstream performance after the middle layers of the model.
@@ 4.1 A Critical Window
In each dataset, correct answers crucially rely on the contextualization of a subject entity (i.e., a polysemous word, a profession, or a country). We begin our investigation by analyzing attention to the subject entities at each layer by all subsequent tokens (which comprise the question). Specifically, for each layer $l$, we compute the mean attention paid to the (final token of the) subject entity over all heads by all subsequent tokens, which we call the attention mass:
@@
$\text{diff}_{y,l}=$ $\frac{\sum_{h_{l}\in D_{y}^{+}}\text{log\_odds}(h_{l}))}{|D_{y}^{+}|}$ (2)
$-\frac{\sum_{h_{l}\in D_{y}^{-}}\text{log\_odds}(h_{l}))}{|D_{y}^{-}|}$
This value quantifies when the logit differences of a successful model generation differ from the logit differences of a failure, while controlling for intrinsic biases toward producing affirmative or negative answers that occur throughout the layers of a model (see § C). From Fig. 4, we see that there is no difference between success and failures until the middle layers of the network. After this point, there is clear signal indicating the answer that the model will ultimately generate. Thus, we might pin down the end of the critical window to approximately layer 20, or shortly thereafter. See § C for disaggregated details, which further support the conclusions presented here, and additional analysis.
@@
Figure 5: Attention Ablation Results. Ablating cue tokens and distractor tokens both have the intended impact on performance – ablating cue tokens drops performance and ablating distractors increases performance. Notably, the interventions only impact model performance in the first half of layers.
Given the qualitative results obtained by analyzing attention and logit lens, we should expect that contextualization should only impact model performance in early and middle layers, before the question tokens finish reading from the subject token(s). We perform causal interventions on the attention maps produced throughout the transformer layers to test this. Specifically, we perform attention ablations: zeroing out the attention map entry that defines how much attention one token pays to another and renormalizing the resulting values. This effectively removes the influence of one token on another within the layers in which we apply this intervention. Formally, let $T_{E}$ be the set of tokens to edit and $T_{A}$ be the set of tokens to ablate, such that $T_{E}$ is the complement of $T_{A}$. Attention ablations are defined as follows: $\forall_{e\in T_{E},a\in T_{A}}A_{e,a}\leftarrow 0$. We use the transformerlens package to implement this intervention (Nanda and Bloom, 2022). We perform two types of attention ablations: distractor ablation, where $T_{A}$ corresponds to the tokens comprising distractor sentences; and cue ablation, where $T_{A}$ corresponds to tokens comprising the cue. Distractor ablation mimics the no distractors setting—up to a difference in positional embeddings—over the layers in which it is applied and should boost performance. Cue ablation should leave the model unable to properly contextualize the subject entity, harming accuracy.
We perform distractor ablation and cue ablation in blocks of 5 layers and analyze the impact of attention ablation over the course of the model. Our main Gemma-2 results are presented in Fig. 5, and a replication with another LLM is presented in § G. In all cases, we find that distractor ablation reliably increases performance in early layers, and cue ablation reliably decreases performance in early layers. Furthermore, contextualization interventions stop impacting model performance at approximately halfway through the model layers—almost precisely when logit lens reveals contextualization of the final question token.
4.2 Subject Entity Contextualization
Next, we verify that model failures are caused by unresolved contextualization of the subject entity in early layers of the model (i.e., before the end of the critical window). We support this by (1) analyzing the content of subject entity representations over layers and finding that distractors deteriorate contextualization, (2) performing causal patching analyses which show that contextualization of the subject entity mediates model performance and that well-contextualized representations emerge in late layers of the model.
@@ 4.2 Subject Entity Contextualization
First, we qualitatively investigate how the content of subject entity representations evolves over layers of processing. To do so, we employ an open-ended patchscope (Ghandeharioun et al., 2024a) inspired by the SelfIE technique (Chen et al., 2024). This technique formulates a patchscope where the target prompt, $T$, is an interpretation prompt—a prompt designed to encourage the model to describe the hidden representation. Source representations are extracted from $i$, where $i$ is the position of the last token of the subject entity, and $l$ ranges over every even layer (i.e., $l\in[0,2,4...,N]$) in an LLM, $\mathcal{M}$. The source representations are patched into a target defined by $T=$Tell me about X X X”, $i^{*}=$ the index of the X tokens, and $l^{*}=3$ (following Chen et al., 2024). We then condition the model to respond with “Sure! In this context, the word refers to” and allow the model to freely generate 15 tokens. This process ideally results in a natural language description of the content of the hidden representation.
We employ this technique to investigate how the word sense of the subject entities in our polysemous word dataset evolve over the layers of the model. Following prior work (Ghandeharioun et al., 2024b), we use another LLM [^3] PaLM2 text-unicorn@001 (Anil et al., 2023) to autoscore these generations as being indicative of one sense or another. See § D for details and § E for example interpretations. We present the results of this analysis in Fig. 6, showing cumulative accuracies over layers. A question pair is marked correct at layer $m$ if the model has properly contextualized the intermediate representation in at least one layer [0, $m$]. In Fig. 6 (left), we compare question pairs with and without distractors. We find that the presence of distractors delays proper contextualization of intermediate representations, shifting the cumulative accuracy curve for question pairs with distractors to the right. This delay may limit the number of layers within the critical window where the subject tokens are well-contextualized. Furthermore, in Fig. 6 (right), we compare distractor question pairs that the model ultimately answers correctly with those that the model ultimately answers incorrectly. We find that cumulative accuracy is lower overall in failure cases, indicating that subject entities are sometimes not properly contextualized at all during model failures.
@@
At minimum, the LLM Race Conditions Hypothesis implies that improperly contextualized subject-entity representations somewhere in the model are causing incorrect responses. Behaviorally, we know that the model achieves ceiling performance without distractors, implying that the model contains a properly contextualized subject-entity representation. Can we boost performance in the presence of distractors by injecting subject-entity representations generated by models when they are faced with data without distractors? To test this, we perform a patchscopes intervention. Source representations are specified where $S$ corresponds to a no-distractors prompt, $i$ corresponds to the position of the last token of the subject entity in $S$, $\mathcal{M}$ corresponds to an LLM with $N$ layers, and $l\in[0,2,4,6,...,M]$, where $M=N/2$. Target representations are specified where $T$ corresponds to a prompt with distractors, $i^{*}$ refers to the position of the last token of the subject entity in $T$, and $l^{*}=l$. We refer to this intervention as cross-patching. We search over layers for each question pair separately. If any source layer-target layer combination results in correct generations for both questions, we consider the question pair correct. From Fig. 7, one can see that cross-patching reliably boosts performance across all datasets.
@@
The LLM Race Conditions Hypothesis requires that subject entity contextualization must happen either before or during the critical window of question contextualization. Otherwise, the properly contextualized subject entity would not be integrated in the final prediction. To test this, we employ a backpatching intervention described in Biran et al. (2024): source representations are extracted from a prompt, $S$, at some position $i$ and layer $l$. Target representations are defined by $T=S$, $i^{*}=i$ and $l^{*}<l$. In our case, we let $l^{*}\in[0,2,4,...,M]$ and $l\in[M,M+2,M+4,...,N]$, where the LLM $\mathcal{M}$ has $L$ layers and $M=L/2$. For each layer in the target set, we attempt to patch in a representation from each layer in the source set. In Fig. 7, we find that this reliably increases performance across all datasets.
@@
In the previous experiment, the backpatched representations undergo further processing through the MLP and attention mechanisms after the target layer, $l^{*}$. Perhaps the performance advantages conferred by backpatching are a result of this extra computation, rather than the relative ordering of question contextualization and subject entity contextualization? To address this, we rerun the same experiment and source-target layer search, except we patch the source representation into every layer in between the source and target. Formally, if $l^{\prime}$ indicates the set of target layers, $l^{\prime}=[l^{*},l^{*}+1,...,l]$. This effectively freezes the representation, rendering it insensitive to MLP and attention mechanisms throughout $l^{\prime}$. From Fig. 7, we see that this setting produces similar results as the standard backpatching setting. Thus, the ordering of contextualization, rather than extra computation, engenders higher performance. See § F for further analysis, including results from every source-target layer combination.